In how many ways can a group of 24 students be organized into groups of 2 if student number 3 and student number 22 must always be one of the pairs? How can this situation be expressed mathematically?
Answer by stanbon(75887)

In how many ways can a group of 24 students be organized into groups of 2 if student number 3 and student number 22 must always be one of the pairs? How can this situation be expressed mathematically?
----
Let those 2 form one of the pairs.
# of other pairs = 22C2 = (22*21)/2 = 11*21 = 231 other pairs
-----
Total number of pairs = 1 + 22C2
-----------------------------------
cheers,
Stan H.
-------------